OpenStudy (anonymous):

how do you solve? square root of 3 divided by square root of 5 ?

OpenStudy (phi):

You don't solve these, but you do put them into a "standard form". People do not like square roots in the denominator, so you should "get rid of it" by multiplying by sqrt(5)/sqrt(5) \[ \frac{\sqrt{3}}{\sqrt{5}} \cdot \frac{\sqrt{5}}{\sqrt{5}}= \frac{\sqrt{15}}{5} \]

OpenStudy (anonymous):

thank you , so you canot simplify it further?

OpenStudy (phi):

If 15 could be factored into a perfect square you could do more, but 3*5 is all you can do with 15, and neither is a square. if you had sqrt(25) you could rewrite it as sqrt(5*5) = 5 if you had sqrt(18) you could rewrite it as sqrt(3*3*2)= 3sqrt(2) but 15? forget about it!